# RLIE: Rule Generation with Logistic Regression, Iterative Refinement, and Evaluation for Large Language Models

### v1

@@ 3.2 Logistic Regression
Feature Construction. For any iteration $t$, given a rule set $\mathcal{H}^{(t)}=\{h_{1}^{(t)},\ldots,h^{(t)}_{m^{(t)}}\}$ where $m^{(t)}\leq H$, we define a mapping $\Phi^{(t)}$ from an input sample to a vector of rule judgments
$\Phi^{(t)}:\ \mathcal{X}\to\{-1,0,+1\}^{m^{(t)}},\qquad\Phi^{(t)}(x_{i})=\mathbf{z}_{i}^{(t)}=\begin{bmatrix}z_{i,1}^{(t)},\ldots,z_{i,m^{(t)}}^{(t)}\end{bmatrix}^{\top}$.
Weight Learning. The probability of a sample $x_{i}$ being classified as positive is modeled by a logistic regression function, which is defined as below
$p^{(t)}(x_{i};\theta^{(t)})=\sigma\!\big((\Phi^{(t)}(x_{i}))^{\top}\bm{\beta}^{(t)}+b^{(t)}\big),\qquad\sigma(u)=(1+e^{-u})^{-1}$
where $\sigma$ is the sigmoid function and $\theta^{(t)}=(\bm{\beta}^{(t)},b^{(t)})$ are the model parameters. We learn these parameters by minimizing the cross-entropy loss on $\mathcal{S}_{\mathrm{tr}}$ with Elastic Net (Zou & Hastie, 2005) regularization:
$\bm{\beta}^{(t)},b^{(t)}=\underset{{\bm{\beta}^{(t)},b^{(t)}}}{\arg\min}\ \frac{1}{|\mathcal{S}_{\mathrm{tr}}|}\!\sum_{x_{i}\in\mathcal{S}_{\mathrm{tr}}}\mathcal{L}(y_{i},p^{(t)}(x_{i};\theta^{(t)}))+\lambda\!\left(\alpha\|\bm{\beta}^{(t)}\|_{1}+\tfrac{1-\alpha}{2}\|\bm{\beta}^{(t)}\|_{2}^{2}\right),$
where $\mathcal{L}$ is the binary cross-entropy loss. The $L_{1}$ penalty encourages sparsity (rule selection), while the $L_{2}$ penalty enhances robustness. The regularization hyperparameters $(\lambda,\alpha)$ are selected via stratified K-fold cross-validation on $\mathcal{S}_{\mathrm{val}}$. The final parameters $\hat{\theta}^{(t)}=(\hat{\bm{\beta}}^{(t)},\hat{b}^{(t)})$ are obtained by refitting the model on the entire $\mathcal{S}_{\mathrm{tr}}$ with the selected hyperparameters.
Label Prediction. After getting rule weights, the prediction for a sample $x_{i}$ can be calculated as
$\hat{y}_{i}^{(t)}=\mathbb{I}\big[p^{(t)}(x_{i};\hat{\theta}^{(t)})\geq\tau\big]\in\{0,1\},\qquad\text{with }\tau=0.5.$
@@ 3.3 Iterative Refinement
In the first iteration, $\mathcal{H}^{(1)}$ and $\hat{\theta}^{(1)}$ are obtained on randomly selected samples. In the subsequent iterations, given $\mathcal{H}^{(t)}$ and $\hat{\theta}^{(t)}$ with $t\geq 1$, to generate $\mathcal{H}^{(t+1)}$ and $\hat{\theta}^{(t+1)}$, instead of relying on randomly selected samples, we can enhance the rule generation by targeting hard examples.
1. Hard Example Selection. For each example $x_{i}\in\mathcal{S}_{\mathrm{tr}}$, we compute its predicted probability $\hat{p}_{i}^{(t)}=p^{(t)}(x_{i};\hat{\theta}^{(t)})$ and define its prediction error as $d_{i}=\|\hat{p}_{i}^{(t)}-y_{i}\|$. We identify the top-$k$ samples with the highest error $I_{\mathrm{hard}}^{(t)}=\mathrm{TopK}\left(\{(i,d_{i})\}^{\|\mathcal{S}_{\mathrm{tr}}\|}_{i=1},k\right)$.
2. New Rule Generation. The selected hard examples $I^{(t)}_{\mathrm{hard}}$ are presented to the LLM along with the current rule set $\mathcal{H}^{(t)}$. The LLM is prompted to reflect on the errors and either revise existing rules or generate new ones that cover novel perspectives. The $h$ newly generated rules are filtered for coverage, resulting in the set $\mathcal{H}_{\mathrm{new}}^{(t+1)}$.
3. Rule Set Update. The new rules are merged with the existing set: $\mathcal{H}_{\mathrm{tmp}}^{(t+1)}=\mathcal{H}^{(t)}\cup\mathcal{H}_{\mathrm{new}}^{(t+1)}$. If the size of this temporary set does not exceed the capacity $H$, it can be directly used as $\mathcal{H}^{(t+1)}$, otherwise we have to prune the rules by ranking them based on their individual accuracy on the validation set, retaining only the top $H$ rules to form $\mathcal{H}^{(t+1)}$.
4. Parameter Update and Termination. With the updated rule set $\mathcal{H}^{(t+1)}$, we calculate $\hat{\theta}^{(t+1)}$ in the same way as discussed in Section 3.2, after which we can monitor the overall performance on $\mathcal{S}_{\mathrm{val}}$. Rules will be refined iteratively, with each iteration trying to improve the performance on hard examples of previous version of rule sets. The process terminates if the overall performance on $\mathcal{S}_{\mathrm{val}}$ fails to improve by a margin $\delta$ for $p$ consecutive iterations or if the maximum number of iterations $R_{\mathrm{max}}$ is reached. The final model $(\mathcal{H}^{\star},\hat{\theta}^{\star})$ corresponds to the checkpoint with the best validation performance.
@@ 3.4 Evaluation
We systematically compare four inference strategies on the test set $\mathcal{S}_{\mathrm{te}}$, categorized into two main approaches: direct inference using the linear combiner and LLM-augmented inference with varying levels of information.
(E1) Linear-only. The final prediction is made directly by the logistic regression model, with $\hat{y}_{i}^{\mathrm{final}}=\mathbb{I}\big[p(x_{i};\hat{\theta}^{\star})\geq 0.5]$, where $p(\cdot;\hat{\theta}^{\star})$ is the learnt probabilistic combiner. This serves as a classic baseline for weighted rule sets.
(E2) LLM + Rules. The LLM is provided with only the texts of the rules in $\mathcal{H}^{\star}$ and the input $x_{i}$, and is tasked with making a prediction where $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathcal{H}^{\star})$. This evaluates the LLM’s intrinsic ability to aggregate rule-based evidence, mirroring common practices in existing work.
(E3) LLM + Rules + Weights. In addition to the rule texts, the LLM is also provided with the learnt weights $\hat{\theta}^{\star}$ and the final prediction becomes $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathcal{H}^{\star},\hat{\theta}^{\star})$. This tests whether the LLM can leverage explicit probabilistic signals to guide its reasoning, reflecting a “probabilistic internalization” of rule importance.
(E4) LLM + Rules + Weights + Linear Prediction. The LLM receives all information from generated rules and corresponding weights, plus the prediction $\hat{y}_{i}$ from the linear-only model as a reference, so the prediction becomes $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathcal{H}^{\star},\hat{\theta}^{\star},\hat{y}_{i})$. This strategy assesses whether a potentially imperfect but calibrated rule-based system can provide a valuable external signal to enhance the LLM’s reasoning in complex scenarios where linear assumption may not fully hold.
@@ 4.1 Dataset
To evaluate our method, we select six real-world tasks from the HypoBench(Liu et al., 2025) benchmark, including Deception Detection (Review), Persuasive Argument Prediction (Dreddit), News Headline Engagement (Headlines), Paper Citations (Citations), AI-generated Content Detection (LLM Detect) and Retweets (Retweets). Each of the task can be viewed as a binary classification task, with text information as inputs and binary classification results as output (e.g. given a review, we need to evaluate whether it is deceptive or not). Details of each dataset are in Appendix A.3.
@@ 4.2 Baseline
We implemented and compared RLIE against several representative and competitive methods within a unified framework to comprehensively evaluate its performance and robustness. (1) Zero-shot Inference: The LLM directly predicts the output for a given sample based only on the task description, without access to any examples or rules. (2) Zero-shot Generation: The LLM generates a set of candidate rules based solely on the task description. The single best-performing rule, as determined on the validation set, is then used for inference on the test set. (3) IO Refinement (Qiu et al., 2023): This method follows a “propose-select-refine” loop. In each round, the LLM generates a batch of candidate hypotheses, the best one is selected based on its performance on a validation or error set, and it is then used as a seed for the next round of refinement. This process is repeated for a fixed number of iterations. (4) HypoGeniC (Zhou et al., 2024): This approach maintains and dynamically expands a library of hypotheses. For each sample, the top-$k$ hypotheses are selected to make a prediction. Mismatched samples are added to an error set, which triggers the generation of new hypotheses to improve coverage. The hypothesis library is continuously updated based on a reward signal.
@@ 4.3 Experimental Details
In our experiments, each dataset was partitioned into fixed-size training, validation, and test sets of 200, 200, and 300 samples, respectively. All methods used the same data splits and a fixed random seed to ensure reproducibility. We report Accuracy and Macro-F1 score. Each experiment was repeated at least three times, and we report the mean and standard deviation of the results. All experiments involving LLMs utilized gpt-4o-mini with the temperature set to $1\times 10^{-5}$ to ensure deterministic outputs. Prompt details can be found in Appendix C. As for hyperparameters, the capacity of the rule set was set to $H=10$. In each iteration, we sampled $k=20$ hard examples and generated $h=5$ new rules. The minimum coverage threshold for rules was set to $\gamma=0.2$. The implementations of IO Refinement and HypoGeniC followed the default hyperparameter settings of the original papers. Further details are provided in Appendix A.
@@ 5 Results
This section presents empirical evaluation of our framework. First, we compare the overall performance of RLIE against competitive baselines. Second, we conduct an ablation study on different inference strategies to identify the most effective way to utilize the learned probabilistic rules.
Table 1: Overall performance of different frameworks evaluated in Accuracy and F1 Score ($mean\pm std$, $\times 10^{-2}$). The best-performing results are shown in bold, while the second best results are highlighted in underline. The performance of RLIE is evaluated with the Linear Only strategy.
Method Reviews Dreaddit Headlines Citations LLM Detect Retweets
Accuracy
Zero-shot Inference 55.11±0.18 62.00±0.88 63.22±2.52 56.25±0.00 56.11±0.38 65.56±2.34
Zero-shot Generation 63.89±5.50 64.33±2.19 53.67±1.20 58.33±3.61 52.44±1.07 57.67±2.40
IO Refinement 64.11±5.52 74.89±3.47 59.67±3.51 48.96±1.80 60.33±17.32 62.44±0.69
HypoGeniC 58.00±6.35 65.67±0.00 60.78±3.27 47.92±7.22 54.56±4.43 65.56±2.12
RLIE (Ours) 71.00±0.67 69.11±1.17 65.56±1.50 59.38±3.12 58.67±2.65 65.78±1.71
F1 Score
Zero-shot Inference 50.36±2.30 55.67±2.30 62.70±2.53 36.00±0.00 46.77±0.68 64.69±2.20
Zero-shot Generation 63.41±6.09 59.17±2.84 53.10±0.66 42.56±7.04 39.68±1.58 57.64±2.42
IO Refinement 62.03±8.45 73.98±3.59 59.21±3.23 45.66±2.19 54.00±22.42 62.30±0.66
HypoGeniC 56.82±6.29 60.57±0.61 59.56±3.82 41.96±3.40 44.17±8.75 65.49±2.17
RLIE (Ours) 70.81±0.59 65.94±1.25 65.38±1.44 49.08±2.84 52.10±4.51 65.45±1.41
@@ 5.1 Main Results
The empirical results, summarized in Table 1, demonstrate the effectiveness and robustness of RLIE compared to several strong baselines across six real-world datasets. For this comparison, RLIE’s performance is evaluated using its direct inference strategy (Linear-only), where the learned logistic regression model acts as the classifier. While some baseline methods may attain the top performance on one or two individual datasets, RLIE consistently ranks within the top two in terms of both Accuracy and F1 Score. This demonstrates that the RLIE framework can be generalized to diverse data distributions and is effective at capturing underlying data patterns. As shown in the results table, some methods exhibit high variance and instability despite their strong performance on certain tasks (e.g., IO Refinement). In contrast, our method achieves high performance while maintaining stability, underscoring its robustness.
A horizontal comparison of the different methods reveals several interesting insights. It is noteworthy that Zero-shot Inference outperforms many of the more complex baselines in several scenarios. This may be attributed to the extensive knowledge acquired by the LLM during pre-training, which is sufficient for solving certain tasks. However, this internal inference process is inherently unexplainable, and it is difficult to rule out the influence of potential data leakage from the pre-training corpus. This observation highlights both the challenge and the necessity of using rule-based learning to achieve interpretable reasoning. HypoGeniC learns a set of rules independently and fails to account for the interactions between them, resulting in suboptimal performance. By iteratively refining a single rule, IO Refinement achieves respectable results, which validates the value of providing annotated examples and employing an iterative optimization process. It is worth mentioning that in some cases, IO Refinement outperforms RLIE. This could be because the strategy of generating only a single rule forces it to be more generalizable, allowing it to better handle unseen scenarios. However, this single-rule approach also limits the method’s expressiveness, leading to poor performance in cases that require a more complex rule set. Our RLIE framework strikes a balance: by continuously learning from hard examples, it enhances the generalizability of the learned rule set, while the strategy of maintaining multiple rules significantly boosts its expressiveness. This allows our method to achieve excellent and robust performance with low variance across diverse datasets.
Table 2: Performance of different inference strategies valuated in Accuracy and F1 Score ($mean\pm std$, $\times 10^{-2}$). The best-performing results are shown in bold, while the second best results are highlighted in underline.
Inference Strategy Reviews Dreaddit Headlines Citations LLM Detect Retweets
Accuracy
Regression Only 71.00±0.67 69.11±1.17 65.56±1.50 59.38±3.12 58.67±2.65 65.78±1.71
LLM Evaluation 65.22±2.67 64.56±0.51 61.33±3.06 60.42±1.80 55.22±2.12 61.00±3.06
LLM + Linear Regression 64.67±2.33 67.56±2.50 61.89±3.02 60.42±1.80 54.11±2.22 62.22±3.56
LLM + Linear + Label 64.67±2.60 69.00±1.33 65.33±1.86 59.38±3.12 56.67±3.51 60.22±3.56
F1 Score
Regression Only 70.81±0.59 65.94±1.25 65.38±1.44 49.08±2.84 52.10±4.51 65.45±1.41
LLM Evaluation 64.52±3.19 59.57±0.94 60.88±3.18 48.62±1.19 46.82±5.32 60.33±3.13
LLM + Linear Regression 63.87±3.05 63.73±3.72 61.71±2.93 47.11±3.38 44.76±4.62 59.68±4.45
LLM + Linear + Label 63.80±2.87 65.79±1.51 65.16±1.78 49.08±2.84 48.70±6.05 57.38±4.84

### v2

@@ 3.2 Logistic Regression
Unlike neural networks which are sometimes overconfident(Guo et al., 2017), our logistic combiner minimizes log-loss, inherently promoting calibration.
Feature Construction. At iteration $t$, given a rule set $\mathcal{H}^{(t)}=\{h_{1}^{(t)},\ldots,h^{(t)}_{m^{(t)}}\}$ where $m^{(t)}\leq H$, we define a mapping $\Phi^{(t)}$ from an input sample to a vector of rule application results $\Phi^{(t)}:\mathcal{X}\to\{-1,0,1\}^{m^{(t)}},\quad\Phi^{(t)}(x_{i})=\mathbf{z}_{i}^{(t)}=[z_{i,j}^{(t)}]_{j=1}^{m^{(t)}}.$
Weight Learning. We model the positive-class probability using a regularized logistic regression combiner $p^{(t)}(x_{i};\theta^{(t)})=\sigma\!\big((\Phi^{(t)}(x_{i}))^{\top}\boldsymbol{\beta}^{(t)}+b^{(t)}\big)$ where $\sigma$ is the sigmoid function and $\theta^{(t)}=(\boldsymbol{\beta}^{(t)},b^{(t)})$. We learn these parameters by minimizing the cross-entropy (CE) loss on $\mathcal{S}_{\mathrm{train}}$ with Elastic Net (Zou and Hastie, 2005) regularization:
$\begin{split}\boldsymbol{\beta}^{(t)},\,b^{(t)}=\mathop{\arg\min}_{\boldsymbol{\beta}^{(t)},\,b^{(t)}}\Bigg[&\frac{1}{|\mathcal{S}_{\mathrm{train}}|}\sum_{x_{i}\in\mathcal{S}_{\mathrm{train}}}\mathcal{L}\!\left(y_{i},\,p^{(t)}(x_{i};\theta^{(t)})\right)\\ &+\lambda\left(\alpha\|\boldsymbol{\beta}^{(t)}\|_{1}+\frac{1-\alpha}{2}\|\boldsymbol{\beta}^{(t)}\|_{2}^{2}\right)\Bigg]\end{split}$
where $\mathcal{L}$ denotes binary cross-entropy. The $L_{1}$ term promotes sparsity (rule selection), while the $L_{2}$ term improves robustness. We select $(\lambda,\alpha)$ on $\mathcal{S}_{\mathrm{val}}$ via stratified $K$-fold cross-validation, and refit on the full $\mathcal{S}_{\mathrm{train}}$ to obtain $\hat{\theta}^{(t)}$.
Label Prediction. Given $\hat{\theta}^{(t)}$, we predict sample $x_{i}$ with $\hat{y}_{i}^{(t)}=\mathbb{I}\big[p^{(t)}(x_{i};\hat{\theta}^{(t)})\geq 0.5\big]\in\{0,1\}.$
@@ 3.3 Iterative Refinement
Although the initial rule set $\mathcal{H}^{(1)}$ is derived from a random subset of $\mathcal{S}_{\mathrm{train}}$, we subsequently refine it by targeting hard examples under the current model $(\mathcal{H}^{(t)},\hat{\theta}^{(t)})$ for $t\geq 1$.
1. Hard Example Selection. For each $(x_{i},y_{i})\in\mathcal{S}_{\mathrm{train}}$, we compute the predicted probability $\hat{p}_{i}^{(t)}=p^{(t)}(x_{i};\hat{\theta}^{(t)})$ and define the prediction error $d_{i}=|\hat{p}_{i}^{(t)}-y_{i}|$. We then select the indices of the top-$k$ hardest examples: $\mathcal{I}_{\mathrm{hard}}^{(t)}=\mathrm{TopK}\!\left(\{d_{i}\}_{i=1}^{|\mathcal{S}_{\mathrm{train}}|},\,k\right).$
2. New Rule Generation. We present the hard examples $\mathcal{I}_{\mathrm{hard}}^{(t)}$ along with the current rule set $\mathcal{H}^{(t)}$ to the LLM, prompting it to revise existing rules or propose $h$ new rules (see Figure 6). After applying the same coverage-based filtering as in Section 3.1, we obtain $\mathcal{H}_{\mathrm{new}}^{(t+1)}$.
3. Rule Set Update. We merge new rules with the current set: $\mathcal{H}_{\mathrm{tmp}}^{(t+1)}=\mathcal{H}^{(t)}\cup\mathcal{H}_{\mathrm{new}}^{(t+1)}$. If $|\mathcal{H}_{\mathrm{tmp}}^{(t+1)}|\leq H$, we set $\mathcal{H}^{(t+1)}=\mathcal{H}_{\mathrm{tmp}}^{(t+1)}$. Otherwise, we prune rules by ranking them according to their F1 scores on $\mathcal{S}_{\mathrm{val}}$ calculated on non-abstaining samples and retaining the top $H$ rules.
4. Parameter Update and Termination. Given $\mathcal{H}^{(t+1)}$, we refit the logistic regression combiner as described in Section 3.2 to obtain $\hat{\theta}^{(t+1)}$. The refinement terminates if validation performance fails to improve by at least $\delta$ for $p$ consecutive iterations, or if the maximum iteration count $R_{\mathrm{max}}$ is reached. We return $(\mathcal{H}^{\star},\hat{\theta}^{\star})$ from the checkpoint with the best validation performance.
@@ 3.4 Evaluation
Upon obtaining the weighted rule sets, we employ various inference strategies to verify their quality and identify the optimal deployment protocol. These strategies fall into two paradigms: direct inference using the learned logistic combiner, and LLM-based inference where the prompt is progressively augmented with information from the learned system. Let $\mathrm{LLM}(\cdot)$ denote an LLM that outputs the final label. We use $\mathbf{R}$ to denote the rule texts in $\mathcal{H}^{\star}$, $\mathbf{W}$ the learned weights $\hat{\theta}^{\star}$, and $\mathbf{P}$ the combiner’s predicted label $\hat{y}_{i}^{\mathrm{comb}}=\mathbb{I}\!\left[p(x_{i};\hat{\theta}^{\star})\geq 0.5\right]$.
(E1) Combiner. We set $\hat{y}_{i}^{\mathrm{final}}=\hat{y}_{i}^{\mathrm{comb}}$, directly utilizing the combiner’s output as the final prediction without further LLM prompting at inference time.
(E2) LLM(R). The LLM is provided with the input $x_{i}$ and rule texts $\mathbf{R}$: $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathbf{R})$. This tests the LLM’s intrinsic ability to reason over a set of natural-language rules absent of explicit weight signals.
(E3) LLM(RW). The LLM is provided with $x_{i}$, $\mathbf{R}$, and the learned weights $\mathbf{W}$: $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathbf{R},\mathbf{W})$. This tests whether explicit probabilistic importance cues facilitate the LLM in prioritizing and aggregating rule evidence.
(E4) LLM(RWP). The LLM is provided with $x_{i}$, $\mathbf{R}$, $\mathbf{W}$, and the combiner prediction $\mathbf{P}$: $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathbf{R},\mathbf{W},\mathbf{P})$. This tests whether the combiner’s prediction can serve as a reference signal to stabilize LLM-based aggregation in challenging cases.
@@ 4.1 Dataset
We evaluate RLIE on six real-world tasks, including Reviews (deceptive review detection), Dreaddit (mental stress detection), Headlines (headline engagement prediction), Citations (paper citation impact prediction), LLM Detect (AI-generated text detection), and Retweets (pairwise retweet prediction). All tasks are formulated as binary classification over natural-language inputs (e.g. given a review, we need to evaluate whether it is deceptive or not). Detailed dataset descriptions are provided in Appendis A.3.
Table 1: Overall performance comparison (Accuracy / Macro-F1, $\times 10^{-2}$) on full datasets. RLIE (Ours) uses the Linear-Only inference strategy. The best results among generalizable methods are bolded.
Method Backbone Reviews Dreaddit Headlines Citations LLM Detect Retweets
Zero-shot DeepSeek-V3.2 53.9 / 42.3 64.7 / 60.0 61.2 / 59.6 62.5 / 50.0 82.4 / 82.0 63.5 / 62.6
Few-shot (ICL) DeepSeek-V3.2 65.3 / 65.1 63.8 / 58.5 62.0 / 61.8 60.4 / 49.8 80.4 / 79.6 57.7 / 53.9
Zero-shot Gen DeepSeek-V3.2 65.4 / 64.8 67.2 / 63.9 57.5 / 57.4 46.1 / 47.7 63.1 / 57.6 58.1 / 58.1
IO Refinement DeepSeek-V3.2 65.9 / 65.4 78.5 / 78.1 62.0 / 61.1 54.2 / 51.0 83.6 / 83.4 57.1 / 56.1
HypoGeniC DeepSeek-V3.2 69.1 / 69.3 80.5 / 80.5 59.9 / 60.1 46.9 / 49.3 85.2 / 85.1 61.9 / 61.8
RLIE (Ours) Qwen3-Next-80B 68.3 / 67.8 81.1 / 81.1 61.1 / 60.9 56.3 / 55.0 87.6 / 87.6 61.9 / 61.8
RLIE (Ours) Qwen3-235B 71.5 / 71.4 79.9 / 79.9 60.6 / 60.4 61.5 / 60.6 88.3 / 88.3 66.5 / 66.5
RLIE (Ours) DeepSeek-V3.2 70.9 / 70.7 82.3 / 82.3 67.0 / 67.0 64.6 / 63.0 90.7 / 90.7 65.7 / 65.6
@@ 4.2 Baseline
We compare RLIE with representative LLM-based inference and rule-learning baselines: (1) Zero-shot: direct prediction from task description without labeled demonstrations or rules. (2) Few-shot (ICL): standard in-context learning with a small set of labeled demonstrations sampled from the training split. (3) Zero-shot Gen: generate a pool of candidate rules from task description only, then select the single best rule on the validation set for test-time inference. (4) IO Refinement (Qiu et al., 2023): iteratively propose-select-refine a single rule using an error-driven refinement loop. (5) HypoGeniC (Zhou et al., 2024): maintain and expand a hypothesis library, and perform top-$k$ hypothesis selection for prediction with reward-driven updates.
@@ 4.3 Experimental Details
LLM backbones. To ensure a fair comparison for all methods that generate or refine rules, we use DeepSeek-V3.2 as the default backbone throughout the main experiments. To examine whether our rule-generation framework is model-agnostic, we additionally run RLIE with alternative backbones (Qwen3-Next-80B and Qwen3-235B) while keeping the rest of the pipeline unchanged.
Protocol and metrics. We use the original train/validation/test splits provided by HypoBench for all methods, and report mean Accuracy (ACC) and Macro-F1 on the test set over three runs with different random seeds.
RLIE hyperparameters. We set the rule capacity to $H=10$. In each iteration, we sampled $k=20$ hard examples and generated $h=5$ new rules. We filter candidate rules by a minimum training coverage threshold $\gamma=0.2$. During iterative refinement, we set the max iteration number $R_{\mathrm{max}}$ to be $20$. The sampling temperature is set to be $1\times 10^{-5}$ for stability and reproducibility. The implementations of IO Refinement and HypoGeniC follow the original papers. Further details are provided in Appendix A.
@@ 5 Results
In this section, we analyse the overall performance of RLIE, conducting experiments on different inference strategies to find best way to apply the learnt rules. We also provide computational cost analysis and two case studies.
@@ 5.1 Main Results
To enable a direct comparison with baselines, we evaluate RLIE using the combiner for final prediction throughout this section. The results are summarized in Table 1, showing that RLIE is strongly competitive across all tasks. We further evaluate RLIE with three different backbones and observe a consistent trend: stronger backbones typically yield a higher overall ceiling, while even with a smaller backbone RLIE remains highly competitive on multiple tasks, indicating robust adaptability across different model capacity regimes.
Comparing against interpretable rule-based baselines reveals a clear progression in capability. Zero-shot Gen generates a set of candidate rules and selects the best single rule for inference, which is limited in expressiveness as it relies on a single hypothesis. IO Refinement iteratively improves a single rule, typically leading to stable gains by enhancing rule quality. HypoGeniC extends this to a set of multiple rules, which can provide additional benefits in some cases but does not consistently outperform single-rule methods, suggesting that multi-rule inference requires a more reliable global integration mechanism. Building on these observations, RLIE couples iterative rule expansion with weight learning and global aggregation: it progressively addresses uncovered hard cases while explicitly learning rule contributions and combining them at the global level, thereby leveraging the benefits of rule sets more reliably and achieving consistent improvements across most tasks.
We also consider direct prompting (Zero-shot/Few-shot) in which LLMs directly output judgments without producing explicit rules, thus offering weaker interpretability. Notably, Few-shot does not necessarily improve over Zero-shot, indicating that inducing implicit generalization from a handful of examples is challenging, often failing to capture stable patterns from insufficient samples.